# Incompressible Limit of Compressible Ideal MHD Flows inside a Perfectly Conducting Wall

### v4

@@ Abstract
Keywords: Compressible ideal MHD, Incompressible limit, Perfectly conducting wall. MSC(2020) codes: 35L60, 35Q35, 76M45, 76W05.
@@ 1 Introduction
In this paper, we consider the compressible ideal magnetohydrodynamics(MHD) equations
@@
describing the motion of a compressible conducting fluid in an electro-magnetic field. Here $\Omega:=\mathbb{T}^{d-1}\times(-1,1)$ is the reference domain in $\mathbb{R}^{d}~{}(d=2,3)$ with boundary $\Sigma:=\Sigma_{+}\cup\Sigma_{-}$ and $\Sigma_{\pm}:=\{x_{d}=\pm 1\}$. $\nabla:=(\partial_{x_{1}},\cdots,\partial_{x_{d}})$ is the standard spatial derivative. $D_{t}:=\partial_{t}+u\cdot\nabla$ is the material derivative. The fluid velocity, the magnetic field, the fluid density, the fluid pressure and the entropy are denoted by $u=(u_{1},\cdots,u_{d})$, $B=(B_{1},\cdots,B_{d})$, $\rho$, $p$ and $S$ respectively. Note that the last equation of (1.1) is derived from the equation of total energy and Gibbs relation. See [6, Ch. 4.3] for more details. Throughout this paper, we will use follow the convention of Einstein’s summation, that is, repeated indices represent taking sum over the indices.
@@
These two conditions also guarantee the hyperbolicity of system (1.1). The initial and boundary conditions of system (1.1) are
@@
$\partial_{t}^{j}u_{d}|_{t=0}=0,~{}~{}0\leq j\leq m.$ (1.5)
@@
Thus the compressible MHD system is now reformulated as follows
$\begin{cases}\frac{\partial\mathcal{F}}{\partial p}D_{t}p+\nabla\cdot u=0~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ \rho D_{t}u=B\cdot\nabla B-\nabla P,~{}~{}P:=p+\frac{1}{2}|B|^{2}~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ D_{t}B=B\cdot\nabla u-B(\nabla\cdot u)~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ \nabla\cdot B=0~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ D_{t}S=0~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ p=p(\rho,S),\frac{\partial p}{\partial\rho}>0~{}~{}~{}&\text{in}~{}[0,T]\times\bar{\Omega},\\ u_{d}=B_{d}=0~{}~{}~{}&\text{on}~{}[0,T]\times\Sigma,\\ (u,B,\rho,S)|_{t=0}=(u_{0},B_{0},\rho_{0},S_{0})~{}~{}~{}&\text{on}~{}\{t=0\}\times\Omega.\end{cases}$ (1.7)
1.1 The equation of states and sound speed
This paper is devoted to studying the behavior of the solution of (1.7) as the sound speed goes to infinity, which is known to be the incompressible limit. Physically, the sound speed is defined by $c_{s}:=\sqrt{\frac{\partial p}{\partial\rho}}$. Mathematically, Mathematically, it is convenient to view the sound speed as a parameter $\lambda$. This being said, a typical choice of the equation of states $p_{\lambda}(\rho,S)$ would be the polytropic gas
$\rho_{\lambda}(p/\lambda^{2},S)=\left(\left(1+\frac{p}{\lambda^{2}}\right)e^{-\frac{S}{C_{V}}}\right)^{\frac{1}{\gamma}},\quad\text{and}\,\,\log\left(\rho_{\lambda}(p/\lambda^{2},S)\right)=\gamma^{-1}\log\left(\left(1+\frac{p}{\lambda^{2}}\right)e^{-\frac{S}{C_{V}}}\right).$ (1.9)
Hence, we can view $\mathcal{F}=\log\rho$ as a parametrized family $\{\mathcal{F}_{\varepsilon}(p,S)\}$ as well, where $\varepsilon=\frac{1}{\lambda}$. Indeed, we have
$\frac{\partial\mathcal{F}_{\varepsilon}}{\partial p}=\gamma^{-1}\log\left((1+\varepsilon^{2}p)e^{-\frac{S}{C_{V}}}\right).$ (1.10)
Since we work on the case when the entropy and velocity are both bounded (later we will assume $u,S\in H^{4}(\Omega)$), we again slightly abuse the terminology, and call $\lambda$ the sound speed and call $\varepsilon$ the Mach number and thus $M=O(\varepsilon)$. Furthermore, there exists $C>0$ such that
@@ 1.1 The equation of states and sound speed
Also, we assume
$|\partial_{p}^{s}\mathcal{F}_{\varepsilon}(p,S)|\leq C,\quad|\partial_{p}^{s}\mathcal{F}_{\varepsilon}(p,S)|\leq C|\partial_{p}\mathcal{F}_{\varepsilon}(p,S)|^{s}\leq C\mathcal{F}^{\prime}_{\varepsilon}(p)$ (1.12)
holds for $0\leq s\leq 8$. Thus, when considering the incompressible limit, that is, when $\varepsilon>0$ is sufficiently small, it is more convenient to reformulate the compressible MHD system by replacing $\frac{\partial\mathcal{F}}{\partial p}$ with $\varepsilon^{2}$ as follows
$\begin{cases}\varepsilon^{2}D_{t}p+\nabla\cdot u=0~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ \rho D_{t}u=B\cdot\nabla B-\nabla P,~{}~{}P:=p+\frac{1}{2}|B|^{2}~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ D_{t}B=B\cdot\nabla u-B(\nabla\cdot u)~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ \nabla\cdot B=0~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ D_{t}S=0~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ p=p(\rho,S),\frac{\partial p}{\partial\rho}>0~{}~{}~{}&\text{in}~{}[0,T]\times\bar{\Omega},\\ u_{d}=B_{d}=0~{}~{}~{}&\text{on}~{}[0,T]\times\Sigma,\\ (u,B,\rho,S)|_{t=0}=(u_{0},B_{0},\rho_{0},S_{0})~{}~{}~{}&\text{on}~{}\{t=0\}\times\Omega.\\ \end{cases}$ (1.13)
@@ 1.2 An overview of previous results
For compressible ideal MHD, the incompressible limit in the whole space was studied by Jiang-Ju-Li [10]. However, when the domain has a boundary, the study of incompressible limit for ideal compressible MHD is much more subtle. The first difficulty is that the vorticity estimate for ideal compressible MHD cannot be closed when using standard Sobolev spaces, and thus the method presented in [23, 24] is no longer valid. When the magnetic field is not tangential to the boundary, one can use such transversality to compensate the loss of normal derivative arising in the vorticity analysis. See Yanagisawa [35] for the well-posedness result under the condition $B\times N|_{\partial\Omega}=\mathbf{0}$. As for the singular limits, we refer to Ju-Schochet-Xu [12] for the singular limits when both Mach number and Alfvén number converge to zero and Jiang-Ju-Xu [11] for the low Alfvén number limit of incompressible ideal MHD whose local existence was proved via a low-Mach-number regime.
Unfortunately, when the magnetic field is tangential to the boundary, that is, $B\cdot N|_{\partial\Omega}=0$, Ohno-Shirota [20] proved that the linearized problem near a non-zero magnetic field is ill-posed in standard Sobolev spaces. In this case, one has to use suitable anisotropic Sobolev spaces introduced by Chen [3] to proved the well-posedness. See Yanagisawa-Matsumura [36] and Secchi [26, 28].
Therefore, an essential difficulty in establishing the incompressible limit for ideal MHD with $B\cdot N|_{\partial\Omega}=0$ is the “incompatibility of function spaces for the existence”: compressible ideal MHD flow with the perfectly conducting wall condition may not be well-posed in standard Sobolev spaces, while the corresponding incompressible problem is well-posed (see, for example, Gu-Wang [7]). So far, there is only a partial answer for this incompressible limit problem. Ju and the first author [13] proved the incompressible limit in a suitable closed subspace of standard Sobolev space, introduced by Secchi [27], by adding more restrictive constraints for the boundary value of initial data. Specifically, [27, 13] require the initial data to satisfy
@@ 1.3 The main theorems
*$\partial_{*}^{\alpha}:=\partial_{t}^{\alpha_{0}}(\omega\partial_{3})^{\alpha_{d+1}}\partial_{1}^{\alpha_{1}}\cdots\partial_{d}^{\alpha_{d}},~{}~{}\langle\alpha\rangle:=\sum_{j=0}^{d-1}\alpha_{j}+2\alpha_{d}+\alpha_{d+1},$
and define the space-time anisotropic Sobolev norm $*\|\cdot\|_{m,*}$ to be
@@
First, we establish a local-in-time estimate that is uniform in Mach number $\varepsilon$.
Theorem 1.1 (Uniform estimate in $\varepsilon$).
Let $\varepsilon\in(0,1)$ be fixed. Let $(u_{0},B_{0},\rho_{0},S_{0})\in H_{*}^{8}(\Omega)\times H_{*}^{8}(\Omega)\times H_{*}^{8}(\Omega)\times H_{*}^{8}(\Omega)$ be the initial data of (1.13) satisfying the compatibility conditions (1.5) up to 7-th order and
@@
for some $M>0$ independent of $\varepsilon$. Then there exists $T>0$ depending only on $M$, such that (1.13) admits a unique solution $(p(t),u(t),B(t),S(t))$ verifies the energy estimate
@@
$E_{4}(t)=$ $\sum_{k=0}^{4}\left\|\left(\varepsilon^{(k-1)_{+}}\partial_{t}^{k}u,\varepsilon^{(k-1)_{+}}\partial_{t}^{k}B,\varepsilon^{(k-1)_{+}}\partial_{t}^{k}S,\varepsilon^{k}\partial_{t}^{k}p\right)\right\|^{2}_{4-k}$
$E_{4+l}(t)=$ $\sum_{\begin{subarray}{c}\langle\alpha\rangle=2l,\\ \alpha_{0}<2l\end{subarray}}\sum_{k=0}^{4-l}\left\|\varepsilon^{(k-1)_{+}+2l}\mathcal{T}^{\alpha}\partial_{t}^{k}(u,B,S,p)\right\|_{4-k-l}^{2}$
$+\sum_{k=0}^{4-l}\left\|\varepsilon^{(k-1)_{+}+2l}\partial_{t}^{k+2l}(u,B,S),\varepsilon^{k+2l}\partial_{t}^{k+2l}p\right\|_{4-k-l}^{2},~{}1\leq l\leq 4,$
@@
We note that the norm $\|p\|_{4}^{2}$ in $E_{4}(t)$ defined by (1.18) should be replaced by $\|\varepsilon p\|_{0}^{2}+\|\nabla p\|_{3}^{2}$ because we do not have $L^{2}$ estimates of $p$ without $\varepsilon$ weight. We still write $\|p\|_{4}^{2}$ as above for simplicity of notations.
@@
The energy functional $E(t)$ above is considered as a variant of $*\|\cdot\|_{8,*}$ norm at time $t>0$. For different multi-index $\alpha$, we impose different Mach number weights according to the number of tangential derivatives that appear in *$\partial_{*}^{\alpha}$, such that the energy estimate for the slightly modified $*\|\cdot\|_{8,*}$ norm is uniform in $\varepsilon>0$.
@@
Under the hypothesis of Theorem 1.1, we assume that $(u_{0},B_{0},S_{0})\to(u_{0}^{0},B_{0}^{0},S_{0}^{0})$ in $H^{4}(\Omega)$ as $\varepsilon\to 0$ for $\nabla\cdot u^{0}=\nabla\cdot B^{0}=0$ in $\Omega$ with $u_{03}^{0}=B_{03}^{0}=0$ on $\Sigma$. Then it holds that
@@
where we consider the equation of state (1.9) as $\rho=\rho(\varepsilon^{2}p,S)$, that is, a function of $\varepsilon^{2}p$ and $S$. The function $\pi$ satisfying $\nabla\pi\in C([0,T];H^{3}(\Omega))$ represents the fluid pressure for incompressible MHD system (1.19).
Remark 1.5 (Choice of regularity).
We propose $H^{4}$ regularity for the energy functional because lots of commutator estimates require the bound for ¯($\|\overline{\partial}{}^{2}(u,B)\|_{L^{\infty}}$. Recall that $H^{\frac{d}{2}+\delta}\hookrightarrow L^{\infty}$, so it would be convenient to choose $H^{2+\lceil\frac{d}{2}+\delta\rceil}$ regularity, that is, $H^{4}$ for $d=2,3$. The initial data are set in $H^{8}(\Omega)$ in order to guarantee they satisfy the compatibility conditions up to 7-th order and the uniform-in-$\varepsilon$ bound $E(0)\leq M$. Our results can be easily generalized to the case when $H^{4}$ and $H_{*}^{8}$ are replaced by $H^{m}$ and $H_{*}^{2m}~{}(m\geq 4)$ respectively.

### v6

@@ Abstract
Keywords: Compressible ideal MHD, Incompressible limit, Perfectly conducting wall.
MSC(2020) codes: 35L60, 35Q35, 76M45, 76W05.
@@ 1 Introduction
In this paper, we consider the compressible ideal magnetohydrodynamic(MHD) equations
@@
describing the motion of a compressible conducting fluid in an electro-magnetic field. Here $\Omega:=\mathbb{T}^{d-1}\times(-1,1)$ is the reference domain in $\mathbb{R}^{d}~{}(d=2,3)$ with boundary $\Sigma:=\Sigma_{+}\cup\Sigma_{-}$ and $\Sigma_{\pm}:=\{x_{d}=\pm 1\}$. $\nabla:=(\partial_{x_{1}},\cdots,\partial_{x_{d}})^{\mathrm{T}}$ is the standard spatial derivative. $D_{t}:=\partial_{t}+u\cdot\nabla$ is the material derivative. The fluid velocity, the magnetic field, the fluid density, the fluid pressure and the entropy are denoted by $u=(u_{1},\cdots,u_{d})^{\mathrm{T}}$, $B=(B_{1},\cdots,B_{d})^{\mathrm{T}}$, $\rho$, $p$ and $S$ respectively. Note that the last equation of (1.1) is derived from the equation of total energy and Gibbs relation. See [6, Ch. 4.3] for more details. Throughout this paper, we will use follow the convention of Einstein’s summation, that is, repeated indices represent taking sum over the indices.
@@
These two conditions also guarantee the hyperbolicity of system (1.1).
The initial and boundary conditions of system (1.1) are
@@
$B_{d}|_{\{t=0\}\times\Sigma}=0,~{}~{}\partial_{t}^{j}u_{d}|_{\{t=0\}\times\Sigma}=0,~{}~{}0\leq j\leq m.$ (1.5)
It should be noted that $\partial_{t}^{j}B_{d}|_{\{t=0\}\times\Sigma}=0$ also holds for $1\leq j\leq m$ provided (1.5) holds.
1.1 The equation of state and sound speed
This paper is devoted to studying the behavior of the solution of (1.1) as the sound speed goes to infinity, which is known to be the incompressible limit. Physically, the sound speed is defined by $c_{s}:=\sqrt{\frac{\partial p}{\partial\rho}}$. Mathematically, it is convenient to view the sound speed as a parameter $\lambda$. A typical choice of the equation of states $p_{\lambda}(\rho,S)$ would be the polytropic gas
$\rho_{\lambda}(p,S)=\left(\left(1+\frac{p}{\lambda^{2}}\right)e^{-\frac{S}{C_{V}}}\right)^{\frac{1}{\gamma}},\quad\text{and}\,\,\log\left(\rho_{\lambda}(p,S)\right)=\gamma^{-1}\log\left(\left(1+\frac{p}{\lambda^{2}}\right)e^{-\frac{S}{C_{V}}}\right).$ (1.7)
@@ 1.1 The equation of state and sound speed
Hence, we can view $\mathcal{F}=\log\rho$ as a parametrized family $\{\mathcal{F}_{\varepsilon}(p,S)\}$ as well, where $\varepsilon=\frac{1}{\lambda}$. Since we work on the case when the entropy and velocity are both bounded (later we will prove $\|u,S,\varepsilon p\|_{H^{4}(\Omega)}$ is bounded uniformly in $\varepsilon$), we again slightly abuse the terminology, and call $\lambda$ the sound speed and call $\varepsilon$ the Mach number and thus $M=O(\varepsilon)$. It is easy to see that there exists $C>0$ such that
$\frac{1}{\rho}\frac{\partial\rho}{\partial p}(p,S)\leq C\varepsilon^{2}.$ (1.9)
Furthermore, if $\varepsilon>0$ is sufficiently small, there exists $C>0$ such that
@@
Also, we have for some $C_{k}>0$
$\frac{\partial\mathcal{F}_{\varepsilon}}{\partial p}=\gamma^{-1}\frac{\varepsilon^{2}}{1+\varepsilon^{2}p},~{}~{}\frac{\partial^{k}\mathcal{F}_{\varepsilon}}{\partial p^{k}}=C_{k}\frac{\varepsilon^{2k}}{(1+\varepsilon^{2}p)^{k}},~{}~{}~{}k\in\mathbb{N}^{*},$ (1.11)
and so
$|\partial_{p}^{k}\mathcal{F}_{\varepsilon}(p,S)|\leq C,\quad|\partial_{p}^{k}\mathcal{F}_{\varepsilon}(p,S)|\leq C\partial_{p}\mathcal{F}_{\varepsilon}(p,S)~{}~{}0\leq k\leq 8,~{}k\in\mathbb{N}.$ (1.12)
Writing $\mathcal{F}_{p}:=\frac{\partial\mathcal{F}}{\partial p}$, system (1.1) is reformulated as follows
$\begin{cases}\mathcal{F}_{p}D_{t}p+\nabla\cdot u=0~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ \rho D_{t}u=B\cdot\nabla B-\nabla P,~{}~{}P:=p+\frac{1}{2}|B|^{2}~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ D_{t}B=B\cdot\nabla u-B(\nabla\cdot u)~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ \nabla\cdot B=0~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ D_{t}S=0~{}~{}~{}&\text{in}~{}[0,T]\times\Omega,\\ p=p(\rho,S),\frac{\partial p}{\partial\rho}>0~{}~{}~{}&\text{in}~{}[0,T]\times\bar{\Omega},\\ u_{d}=B_{d}=0~{}~{}~{}&\text{on}~{}[0,T]\times\Sigma,\\ (u,B,\rho,S)|_{t=0}=(u_{0},B_{0},\rho_{0},S_{0})~{}~{}~{}&\text{on}~{}\{t=0\}\times\Omega.\\ \end{cases}$ (1.13)
When considering the incompressible limit, we sometimes replace $\mathcal{F}_{p}$ by $\varepsilon^{2}$ for simplicity of notations.
@@ 1.2 An overview of previous results
For compressible ideal MHD, the incompressible limit in the whole space was studied by Jiang-Ju-Li [10]. However, when the domain has a boundary, the study of incompressible limit for ideal compressible MHD is much more subtle. The first difficulty is that the vorticity estimate for ideal compressible MHD cannot be closed when using standard Sobolev spaces, and thus the method presented in [24, 25] is no longer valid. When the magnetic field is not tangential to the boundary, one can use such transversality to compensate the loss of normal derivative arising in the vorticity analysis. See Yanagisawa [37] for the well-posedness result under the condition $B\times N|_{\Sigma}=\mathbf{0}$. As for the singular limits, we refer to Ju-Schochet-Xu [12] for the singular limits when both Mach number and Alfvén number converge to zero and Jiang-Ju-Xu [11] for the low Alfvén number limit of incompressible ideal MHD whose local existence was proved via a low-Mach-number regime.
Unfortunately, when the magnetic field is tangential to the boundary, that is, $B\cdot N|_{\Sigma}=0$, Ohno-Shirota [21] proved that the linearized problem near a non-zero magnetic field is ill-posed in standard Sobolev spaces. In this case, one has to use suitable anisotropic Sobolev spaces introduced by Chen [3] to proved the well-posedness. See Yanagisawa-Matsumura [38] and Secchi [27, 29].
Therefore, an essential difficulty in establishing the incompressible limit for ideal MHD with $B\cdot N|_{\Sigma}=0$ is the “incompatibility of function spaces for the existence”: compressible ideal MHD flow with the perfectly conducting wall condition may not be well-posed in standard Sobolev spaces, while the corresponding incompressible problem is well-posed (see, for example, Gu-Wang [7]). So far, there is only a partial answer for this incompressible limit problem. Ju and the first author [13] proved the incompressible limit in a suitable closed subspace of standard Sobolev space, introduced by Secchi [28], by adding more restrictive constraints for the boundary value of initial data. Specifically, [28, 13] require the initial data to satisfy
@@ 1.3 The main theorems
$\partial_{*}^{\alpha}:=\partial_{t}^{\alpha_{0}}(\omega\partial_{3})^{\alpha_{d+1}}\partial_{1}^{\alpha_{1}}\cdots\partial_{d}^{\alpha_{d}},~{}~{}\langle\alpha\rangle:=\sum_{j=0}^{d-1}\alpha_{j}+2\alpha_{d}+\alpha_{d+1},$
and define the space-time anisotropic Sobolev norm $\|\cdot\|_{m,*}$ to be
@@
First, we prove the local existence togther with uniform-in-$\varepsilon$ estimates.
Theorem 1.1 (Local well-posedness and uniform-in-$\varepsilon$ estimates).
Let $\varepsilon\in(0,1)$ be fixed. Let $(u_{0},B_{0},\rho_{0},S_{0})\in H^{8}(\Omega)\times H^{8}(\Omega)\times H^{8}(\Omega)\times H^{8}(\Omega)$ be the initial data of (1.13) satisfying the compatibility conditions (1.5) up to 7-th order and
@@
for some $M>0$ independent of $\varepsilon$. Then there exists $T>0$ depending only on $M$, such that (1.13) admits a unique solution $(p(t),u(t),B(t),S(t))$ that verifies the energy estimate
@@
$E_{4+l}(t)=$ $\sum_{\langle\alpha\rangle=2l}\sum_{k=0}^{4-l}\left\|\varepsilon^{(k-1)_{+}+2l}\mathcal{T}^{\alpha}\partial_{t}^{k}\left(u,B,S,\mathcal{F}_{p}^{\frac{(k+\alpha_{0}-l-3)_{+}}{2}}p\right)\right\|_{4-k-l}^{2}~{}~{}0\leq l\leq 4,$
@@
We note that the norm $\|p\|_{4}^{2}$ in $E_{4}(t)$ defined by (1.18) should be replaced by $\|\sqrt{\mathcal{F}_{p}}p\|_{0}^{2}+\|\nabla p\|_{3}^{2}$ because we do not have $L^{2}$ estimates of $p$ without $\varepsilon$ weight. We still write $\|p\|_{4}^{2}$ as above for simplicity of notations.
@@
Remark 1.4 (Weights of Mach number of $p$).
In (1.18), the weight of Mach number of $p$ is slightly different from $(u,B,S)$, but such difference only occurs when $\mathcal{T}^{\alpha}$ are full time derivatives and $k=4-l$. In fact, due to $k\leq 4-l$ and $\alpha_{0}\leq\langle\alpha\rangle=2l$, we know $(k+\alpha_{0}-l-3)_{+}$ is always equal to zero unless $\alpha_{0}=2l$ and $k=4-l$ simultanously hold.
@@
The energy functional $E(t)$ above is considered as a variant of $\|\cdot\|_{8,*}$ norm at time $t>0$. For different multi-index $\alpha$, we impose different Mach number weights according to the number of tangential derivatives that appear in $\partial_{*}^{\alpha}$, such that the energy estimate for the slightly modified $\|\cdot\|_{8,*}$ norm is uniform in $\varepsilon>0$.
Remark 1.6 (Choice of regularity).
We propose $H^{4}$ regularity for the energy functional because lots of commutator estimates require the bound for ¯($\|\overline{\partial}{}^{2}(u,B)\|_{L^{\infty}}$. Recall that $H^{\frac{d}{2}+\delta}\hookrightarrow L^{\infty}$, so it would be convenient to choose $H^{2+\lceil\frac{d}{2}+\delta\rceil}$ regularity, that is, $H^{4}$ for $d=2,3$. It should be noted that previous results about local existence of (1.13), e.g, [27, 29, 31], require the initial data to belong to $H_{*}^{2m}(\Omega)$ for $m\geq 8$ instead of $m\geq 4$. In Appendix A, we establish the local well-posedness of (1.13) for initial data belonging to $H_{*}^{8}(\Omega)$ and satisfying the assumption of Theorem 1.1. The energy functional used to prove the local existence is still $E(t)$. With this new existence theorem and following the same strategies, the results in this paper can be easily generalized to the case when $H^{4}$ and $H_{*}^{8}$ are replaced by $H^{m}$ and $H_{*}^{2m}~{}(m\geq 4)$ respectively.
Remark 1.7 (Compatibility conditions and regularity of initial data).
The initial data are assumed to be $H^{8}(\Omega)$ functions instead of merely $H_{*}^{8}(\Omega)$ functions because we have to guarantee they satisfy the compatibility conditions up to 7-th order and the uniform-in-$\varepsilon$ bound $E(0)\leq M$. For example, the 7-th order compatibility conditions requires $\partial_{t}^{7}u_{3}|_{\{t=0\}\times\Sigma}=0$, but $\partial_{t}^{7}u|_{t=0}$ only belongs to $H_{*}^{1}(\Omega)$ whose trace on the boundary may have no meaning if we choose initial data from $H_{*}^{8}(\Omega)$. On the other hand, if we alternatively choose initial data in $H^{8}(\Omega)$, then at least we have $\partial_{t}^{7}u|_{t=0}\in H^{1}(\Omega)$ whose trace on $\Sigma$ can be defined in $H^{\frac{1}{2}}(\Sigma)$. See also [29, Theorem 2.1], [16, Section 8.3] for the discussion on this issue.
@@
Under the hypothesis of Theorem 1.1, we assume that $(u_{0},B_{0},S_{0})\to(u_{0}^{0},B_{0}^{0},S_{0}^{0})$ in $H^{4}(\Omega)$ as $\varepsilon\to 0$ for $\nabla\cdot u_{0}^{0}=\nabla\cdot B_{0}^{0}=0$ in $\Omega$ with $u_{03}^{0}=B_{03}^{0}=0$ on $\Sigma$. Then it holds that
@@
where we consider the equation of state (1.7) as $\rho=\rho(\varepsilon^{2}p,S)$, that is, a function of $\varepsilon^{2}p$ and $S$.Also, there exists a function $\pi$ such that $\nabla\pi\in C([0,T];H^{3}(\Omega))$ satisfies the incompressible MHD system (1.19) and the convergence result
$\rho^{-1}\nabla p\to\varrho^{-1}\nabla\pi\quad\text{weakly-* in }L^{\infty}([0,T];H^{3}(\Omega))\text{ and strongly in }C([0,T];H^{3-\delta}(\Omega)).$